OpenStudy (anonymous):

Suppose the speed of light in a vacuum were only 25.0 mi/h. Find the length of a bicycle being ridden at a speed of 22.0 mi/h as measured by an observer sitting on a park bench, given that its proper length is 1.78 m.

OpenStudy (anonymous):

\[L'=L\sqrt{1-v^2}\]WHere v is measured as a proportion of the c. \[L'=1.78\sqrt{1-(\frac{22}{25})^2}\]I won't insult you intelligence- you can use a calculator